jocy17:

Find the slope and the y-intercept of the equation 4x + 5y = -9.

Spectrum:

Usually there's not a y in these equations...

jocy17:

well thats what my test is asking lmao

xXChudLord420Xx:

@spectrum wrote:
Usually there's not a y in these equations...
bruh what math class you in, wym

xXChudLord420Xx:

to find slope get y on one side x's multiple will then be the slope

jocy17:

thanks !

Treeko:

Slope: -4/5 y-intercept: (0,--9/5)

JamesEDunbanned:

Dummy

xXChudLord420Xx:

@jamesedunbanned wrote:
Dummy
oh my allah

JamesEDunbanned:

Yes my child

Treeko:

Dummy?

xXChudLord420Xx:

shut up shi'a, more like sh-it'a

JamesEDunbanned:

Yes teacher

JamesEDunbanned:

Teach the uneducated

JamesEDunbanned:

The uncomplete

xXChudLord420Xx:

stop trying to be poetic, heretic

Treeko:

Wizard

Treeko:

HEHEHE wizard.

JamesEDunbanned:

Gizzard

AspenIsBest:

xD

JamesEDunbanned:

Fat ahh

xXChudLord420Xx:

I bet bros an illusionist wizard

xXChudLord420Xx:

abjuration #1

AspenIsBest:

@jamesedunbanned wrote:
Fat ahh
Talking abt yourself again?? Wow

Treeko:

leave me alone hmph. AND IM NOT A WIZARD.

JamesEDunbanned:

Undo yundo you now a frog

Treeko:

ha

xXChudLord420Xx:

@treeko wrote:
leave me alone hmph. AND IM NOT A WIZARD.
bro's specialized in enchantment

AspenIsBest:

Bro just wanted help on math and got a whole ahh argument chain fr xD

Treeko:

LOL

HotPockets14:

truth

khloikamren0117:

4x - 5y = 9. Subtract 4x from both sides of the equation. -5y = -4x + 9. Divide both sides of the equation by -5. y = (-4x + 9) / -5. Simplify. y = (4/5)x - (9/5)

Joe348:

To find the slope and y-intercept of the equation 4x + 5y = -9, we need to rearrange the equation into slope-intercept form (y = mx + b), where m represents the slope and b represents the y-intercept. Let's solve it step by step: 4x + 5y = -9 First, isolate the y term: 5y = -4x - 9 Next, divide both sides of the equation by 5: y = (-4/5)x - 9/5 So, the slope is -4/5 and the y-intercept is -9/5.